# Look Before You Leap: A GUI-Critic-R1 Model for Pre-Operative Error Diagnosis in GUI Automation

### v1

@@ C.1 Public Datasets
Android in the Wild (AITW) rawles2023androidinthewild dataset comprises human demonstrations of device interactions, encompassing both screen captures and actions, alongside corresponding natural language instructions. It includes a substantial collection of 715,000 episodes covering 30,000 unique instructions, across four versions of the Android operating system (v10–13) and eight different device models ranging from the Pixel 2 XL to Pixel 6, each with distinct screen resolutions. The dataset involves complex multi-step tasks necessitating nuanced semantic understanding of both language and visual context.
Android-In-The-Zoo (AITZ) zhang2024android dataset contains 18,643 screen-action pairs together with chain-of-action-thought annotations, spanning over 70 Android apps, coupled with 4× useful annotations compared with action coordinate labels only. Based on the screen episodes from AITW rawles2023androidinthewild , the authors generate candidate answers for the screen descriptions, action thinkings and next action descriptions. These candidates are further validated and refined by human to guarantee alignment with the screenshots.
GUI Odyssey lu2024gui is a comprehensive dataset for training and evaluating cross-app navigation agents. GUI Odyssey comprises 7735 episodes, meticulously curated from 6 different mobile devices such as Pixel Pro and Tablet. It encompasses 6 types of cross-app navigation tasks spanning from general system tool use to media entertainment, requiring navigating through 201 different apps and 1399 app combos from various fields such as video, music, reading
Android Multi-annotation EXpo (AMEX) chai2024amex is a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents. Their capabilities of completing complex tasks by directly interacting with the graphical user interface (GUI) on mobile devices are trained and evaluated with the proposed dataset. AMEX comprises over 104K high-resolution screenshots from 110 popular mobile applications, which are annotated at multiple levels. AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions, each averaging 13 steps with stepwise GUI-action chains.
GUICourse chen2024guicourse . In this paper, the author constructed GUIAct, a GUI navigation dataset in website and android scenarios for enhancing VLMs’ knowledge of GUI systems, including 67k single-step and 15k multi-step action instructions.
@@ C.2 GUI-Critic Dataset Components
C.2.1 GUI-Critic-Train
Figure 7 provides a visualization of the data sources utilized in our training set. For the construction of our training dataset, we predominantly selected data from AMEXchai2024amex due to its extensive repository of high-quality GUI operation data. The raw data encompasses decision-making processes enriched with reasoning, which contributes to the reliability of the annotated correct actions. Detailed methodologies for the generation of incorrect decisions and the formulation of Chain-of-Thought (CoT) critiques are elaborated in the main text (Section 3.2). We preserved a higher proportion of samples containing incorrect decisions to facilitate the model’s acquisition of enhanced error correction competencies. Figure 8 illustrates a data sample with the generated CoT process in GUI-Critic-Train dataset.
Figure 8: A data sample in GUI-Critic-Train dataset.
@@ C.2.2 GUI-Critic-Test
Mobile-Instruction Generalization (GUI-I). Data in GUI-I originates from the AMEX dataset chai2024amex . We ensure that the instructions in GUI-I are distinct from those in GUI-Critic-Train. After human annotation, GUI-I retains 656 GUI screenshots, each accompanied by a user instruction, operational history, and a set of candidate actions.
Mobile-Scenario Generalization (GUI-S). We collect 114 test data from the Odyssey dataset lu2024gui for GUI-I. We select the data with applications different from those in GUI-Critic-Train. Specifically, the novel applications are:
@@
Web-Scenario Generalization (GUI-W). It contains 418 samples of web automation, which are randomly sampled from the GUICourse chen2024guicourse . Although the action space in web environments differs from mobile platforms (e.g., limitations on swipe direction and the incorporation of double-tapping), the underlying operational logic largely remains consistent across both scenarios.

### v2

@@ C.1 Public Datasets
Android in the Wild (AITW) [rawles2023androidinthewild] dataset comprises human demonstrations of device interactions, encompassing both screen captures and actions, alongside corresponding natural language instructions. It includes a substantial collection of 715,000 episodes covering 30,000 unique instructions, across four versions of the Android operating system (v10–13) and eight different device models ranging from the Pixel 2 XL to Pixel 6, each with distinct screen resolutions. The dataset involves complex multi-step tasks necessitating nuanced semantic understanding of both language and visual context.
Android-In-The-Zoo (AITZ) [zhang2024android] dataset contains 18,643 screen-action pairs together with chain-of-action-thought annotations, spanning over 70 Android apps, coupled with 4× useful annotations compared with action coordinate labels only. Based on the screen episodes from AITW [rawles2023androidinthewild], the authors generate candidate answers for the screen descriptions, action thinkings and next action descriptions. These candidates are further validated and refined by human to guarantee alignment with the screenshots.
GUI Odyssey [lu2024gui] is a comprehensive dataset for training and evaluating cross-app navigation agents. GUI Odyssey comprises 7735 episodes, meticulously curated from 6 different mobile devices such as Pixel Pro and Tablet. It encompasses 6 types of cross-app navigation tasks spanning from general system tool use to media entertainment, requiring navigating through 201 different apps and 1399 app combos from various fields such as video, music, reading
Android Multi-annotation EXpo (AMEX) [chai2024amex] is a comprehensive, large-scale dataset designed for generalist mobile GUI-control agents. Their capabilities of completing complex tasks by directly interacting with the graphical user interface (GUI) on mobile devices are trained and evaluated with the proposed dataset. AMEX comprises over 104K high-resolution screenshots from 110 popular mobile applications, which are annotated at multiple levels. AMEX includes three levels of annotations: GUI interactive element grounding, GUI screen and element functionality descriptions, and complex natural language instructions, each averaging 13 steps with stepwise GUI-action chains.
GUICourse [chen2024guicourse]. In this paper, the author constructed GUIAct, a GUI navigation dataset in website and android scenarios for enhancing VLMs’ knowledge of GUI systems, including 67k single-step and 15k multi-step action instructions.
@@ 3. Decision: {Action}
C.2.1 GUI-Critic-Train
Figure 7 provides a visualization of the data sources utilized in our training set. For the construction of our training dataset, we predominantly selected data from AMEX[chai2024amex] due to its extensive repository of high-quality GUI operation data. The raw data encompasses decision-making processes enriched with reasoning, which contributes to the reliability of the annotated correct actions. Detailed methodologies for the generation of incorrect decisions and the formulation of Chain-of-Thought (CoT) critiques are elaborated in the main text (Section 3.2). We preserved a higher proportion of samples containing incorrect decisions to facilitate the model’s acquisition of enhanced error correction competencies. Figure 8 illustrates a data sample with the generated CoT process in GUI-Critic-Train dataset.
Figure 8: A data sample in GUI-Critic-Train dataset.
@@ C.2.2 GUI-Critic-Test
Mobile-Instruction Generalization (GUI-I). Data in GUI-I originates from the AMEX dataset [chai2024amex]. We ensure that the instructions in GUI-I are distinct from those in GUI-Critic-Train. After human annotation, GUI-I retains 656 GUI screenshots, each accompanied by a user instruction, operational history, and a set of candidate actions.
Mobile-Scenario Generalization (GUI-S). We collect 114 test data from the Odyssey dataset [lu2024gui] for GUI-I. We select the data with applications different from those in GUI-Critic-Train. Specifically, the novel applications are:
@@
Web-Scenario Generalization (GUI-W). It contains 418 samples of web automation, which are randomly sampled from the GUICourse [chen2024guicourse]. Although the action space in web environments differs from mobile platforms (e.g., limitations on swipe direction and the incorporation of double-tapping), the underlying operational logic largely remains consistent across both scenarios.